# Mutation-Guided Unit Test Generation with a Large Language Model

### v4

@@ IV-B Results and Analysis
TABLE I: Summary of Results
@@
MutGen HumanEval-Java 98.3% 95.8% 89.5% 11 125.9 899
Leetcode-Java 98.4% 94.8% 89.1% 19 149.4 1629
@@
On HumanEval-Java, for both line coverage and branch coverage, MutGen outperforms EvoSuite, EvoSuitemut and Genvanilla. On Leetcode-Java, we have the opposite since EvoSuite achieves the highest scores for coverage, surpassing both LLM-based techniques. For both datasets, however, differences are relatively small, below three percentage points. Although MutGen is primarily designed to improve mutation score, the results also demonstrate that it consistently outperforms Genvanilla in terms of line and branch coverage. This supports previous observations  [28] that improvements in mutation score can also lead to enhanced code coverage.
@@
Regarding the mutation score, PITest generates an average of 11 and 19 mutants per subject on HumanEval-Java and Leetcode-Java, respectively, resulting in a total of 1,144 and 1,900 mutants across the two datasets. MutGen achieves mutation scores of 89.5% and 89.1% on the two datasets. Compared to EvoSuite, this represents large improvements of 28.8% and 51.3%, respectively, corresponding to around 20 and 30 percentage points. Compared to Genvanilla, MutGen also significantly improves the mutation score, though to a lesser extent. Additionally, MutGen also surpasses EvoSuitemut, which specifically aims to maximize the mutation score, to an even larger extent. Indeed, surprisingly, we observe that EvoSuite slightly outperforms EvoSuitemut. This might be due to the latter being developed 10 years ago [30], without any further maintenance. For example, we found that the mutation operators used by EvoSuitemut are limited, resulting in poor performance of the generated test cases when evaluated by PITest. Another tentative explanation is that, to ensure fair comparisons, recall that we set the time budget as the maximum budget used by MutGen, but a large budget may be necessary for EvoSuitemut.
@@
Furthermore, we report the A12 effect size [13] on mutation scores in the last column of Table I. For the comparison between EvoSuite and MutGen, the A12 values are 0.759 and 0.899, indicating that MutGen has a 75.9% and 89.9% probability of achieving higher mutation scores than EvoSuite on HumanEval-Java and Leetcode-Java, respectively. Similarly, when compared with Genvanilla, the A12 values are 0.650 and 0.734, suggesting that MutGen outperforms Genvanilla with 65.0% and 73.4% probabilities on the same datasets. Additionally, based on the results of the paired-sample Wilcoxon signed-rank test, we found that MutGen achieves statistically significant improvements ($\alpha$ = 0.05) in mutation scores compared with EvoSuite, EvoSuitemut, and Genvanilla. However, there is no statistically significant difference in coverage, as most subjects already achieve near 100% coverage across all approaches.
@@
In terms of overhead, MutGen consumes fewer tokens on HumanEval-Java but more tokens on Leetcode-Java. The Leetcode-Java subjects are solutions to medium- and hard-level algorithmic problems, with more complex logic than those in HumanEval-Java. As a result, PITest generates more mutants for them (19 vs. 11), and Llama-3.3 requires more test cases to kill all mutants. Compared with Genvanilla, EvoSuite, and EvoSuitemut, in Leetcode-Java, MutGen consumes more tokens while achieving higher mutation scores as well as better line and branch coverage. In contrast, on HumanEval-Java, MutGen consumes fewer tokens than EvoSuite and EvoSuitemut while still achieving superior performance in both coverage and mutation score. However, differences of that scale in the number of tokens do not have practical implications.
@@
Overall, results suggest that MutGen helps significantly improve mutation scores over alternatives, with large differences, at the expense of slightly lower code coverage when compared to EvoSuite, for the most complex subject (Leetcode-Java). This may be expected as EvoSuite is designed to maximize coverage. For the other subject (HumanEval-Java), the reverse trend is observed. Even when EvoSuite is specifically configured as EvoSuitemut to maximize mutation scores [30], it performs slightly worse than EvoSuite on both datasets. Therefore, to optimize mutation scores and, thus, fault detection, we recommend the use of MutGen, though the impact on coverage is unclear and probably practically negligible.
@@
To evaluate the effectiveness and limitations of MutGen in killing mutants, we analyze the killed ratio across different mutation operators (described in [1]) after running MutGen for four iterations. Tables II and III present the mutant status counts for each mutation operator on HumanEval-Java and Leetcode-Java, respectively. In both tables, Column K indicates the number of killed mutants for each corresponding mutation operator. Column L shows the number of live mutants (covered by not killed), while Column NoCOV represents the number of mutants that are not covered by any test cases. Column Killed Ratio denotes the proportion of killed mutants relative to the total number of mutants.
We observe that, for frequently applied mutation operators, such as Math and NegateConditionals, MutGen is able to kill the corresponding mutants across iterations progressively. After four iterations, MutGen achieves a high killed ratio—for instance, 97.2% for Math on Leetcode-Java. The remaining 2.8% of mutants are mostly uncovered, meaning they are not exercised by any generated test cases.
In contrast, mutation operators that appear less frequently tend to exhibit lower effectiveness. For example, VoidMethodCalls shows a killed ratio of 69.2% on HumanEval-Java, and TrueReturns achieves only 57.7% on Leetcode-Java, representing the weakest effectiveness among all mutation operators in terms of killed ratio.
@@
Based on the mutation scores obtained at the fourth iteration, code summarization appears to have the most significant impact on the effectiveness of MutGen. As discussed in Section II, code comments can sometimes mislead LLMs or omit essential details—such as input formats—that are crucial for guiding test case generation. The fixing stage also shows a notable impact on effectiveness, particularly on the LeetCode-Java dataset, where many generated test cases initially fail to execute successfully. Prior work has observed that LLM-generated test cases often suffer from low execution success rates. After applying fixes, however, a substantial portion of these test cases become valid and contribute to the test objectives. Our evaluation confirms that the fixing stage contributes meaningfully to improved mutation scores.
Lastly, mutation feedback also plays a role in enhancing effectiveness. As revealed in RQ2, while many mutants generated by common mutation operators are relatively easy for LLMs to kill, mutation feedback still provides effectiveness gains by guiding the model more effectively. Future work may further investigate how mutation feedback can be tailored to help LLMs kill harder-to-detect mutants, and which mutation operators benefit the most from such guidance. We leave this exploration as a direction for future research.
@@ IV-C Threats to Validity
The threat to internal validity mainly lies in the correctness of the implementation of MutGen and the experimental scripts. To reduce this threat, we have carefully reviewed our code.
@@
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [31], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance with other LLMs is left as future work.
@@
Randomness. Randomness may impact the effectiveness of EvoSuite, MutGen, and its variant. To mitigate this, and considering computational cost, we ran each tool on every subject three times and reported average results. We also performed statistical significance tests and calculated effect sizes, as detailed in Section IV-B.
Equivalent Mutants. Our approach uses the PITest mutation testing framework, which includes mechanisms to reduce the creation of equivalent mutants. While this mitigates the problem to some extent, the remaining equivalent mutants can reduce the efficiency of MutGen and prevent it from achieving a 100% mutation score.
Oracle Problem. We assume the original code is correct and only address assertion errors that cause Maven build failures. This simplification can lead to the generation of incorrect test oracles and result in false positives.
Data Leakage. While some public code in benchmark datasets may appear in the LLM’s training data (e.g., Llama-3.3), the specific mutants we generate are unlikely to be memorized. Moreover, we compare MutGen with a vanilla prompting baseline, and the observed performance gains suggest that our improvements stem from prompt engineering and strategy design rather than memorization.
@@ V Related Work
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [18, 11, 39], and pointed in recent work from Meta [9, 28], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [18]. Although some prior work [23, 28] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work on guiding LLMs to generate test cases with stronger defect detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [29] and Randoop [53] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [31] has investigated enhancing mutation scores by modifying its fitness functions, which is an objective closely related to our work. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [27, 66, 37]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [61, 71, 72, 73, 45, 57] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [7, 52, 16].
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [33, 46], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [51, 68, 24, 69, 38, 43]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [68, 47, 70, 63], Java [73, 68, 20, 35], and others [68, 47, 74, 21, 67]. Other works target specific testing goals, such as detecting particular bugs [47, 28], API testing [17, 25, 44, 43], or generating test oracles [42, 26, 15]. LLMs have also been integrated with traditional testing techniques, including mocking [49, 76], evolutionary algorithms [62], and program analysis [34, 56, 60, 55]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [18, 28]. As discussed in Section II, high code coverage does not necessarily imply strong fault detection [18]. However, as demonstrated in our experiments (Section IV-B), improving mutation score often leads to increased code coverage as a by-product.

### v5

@@ IV-B Results and Analysis
TABLE I: Summary of Results: Comparison of MutGen and baseline approaches on different datasets. For each approach and dataset, we report the average per-subject line coverage (Line Cov), branch coverage (Branch Cov), mutation score (Mutation Score), number of mutants per subject (#Mutants per Subject), Vargha and Delaney’s A12 effect size (A12 effect size), total runtime in seconds (Time (secs)), and the total number of tokens consumed (#Tokens). “*” denotes cases when MutGen achieves significantly better results than the compared baselines.
@@
MutGen HumanEval-Java 98.3% 95.8% 89.5%* 11 125.9 899
Leetcode-Java 98.4% 94.8% 89.1%* 19 149.4 1629
@@
On HumanEval-Java, for both line coverage and branch coverage, MutGen outperforms EvoSuite, EvoSuitemut and Genvanilla. On Leetcode-Java, we see the opposite: EvoSuite achieves the highest coverage scores, surpassing both LLM-based techniques. For both datasets, however, differences are relatively small, below three percentage points. Although MutGen is primarily designed to improve mutation score, the results also demonstrate that it consistently outperforms Genvanilla in terms of line and branch coverage. This supports previous observations  [foster2025mutation] that improvements in mutation score can also lead to enhanced code coverage.
@@
Regarding the mutation score, PITest generates an average of 11 and 19 mutants per subject on HumanEval-Java and Leetcode-Java, respectively, resulting in a total of 1,144 and 1,900 mutants across the two datasets. MutGen achieves mutation scores of 89.5% and 89.1% on the two datasets. Compared to EvoSuite, this represents large improvements of 28.8% and 51.3%, respectively, corresponding to around 20 and 30 percentage points. Compared to Genvanilla, MutGen also significantly improves the mutation score, though to a lesser extent. Additionally, MutGen also surpasses EvoSuitemut, which specifically aims to maximize the mutation score, to an even larger extent. Indeed, surprisingly, we observe that EvoSuite slightly outperforms EvoSuitemut. This might be due to the latter being developed 10 years ago [fraser2014large] and receiving no further maintenance. For example, we found that the mutation operators used by EvoSuitemut are limited, resulting in poor performance of the generated test cases when evaluated by PITest. Another tentative explanation is that, to ensure fair comparisons, recall that we set the time budget to the maximum used by MutGen, but a larger budget may be necessary for EvoSuitemut.
@@
Furthermore, we report the A12 effect size [arcuri2011practical] on mutation scores in the third-to-last column of Table I. For the comparison between EvoSuite and MutGen, the A12 values are 0.759 and 0.899, indicating that MutGen has a 75.9% and 89.9% probability of achieving higher mutation scores than EvoSuite on HumanEval-Java and Leetcode-Java, respectively. Similarly, when compared with Genvanilla, the A12 values are 0.650 and 0.734, suggesting that MutGen outperforms Genvanilla with 65.0% and 73.4% probabilities on the same datasets. Additionally, based on the results of the paired-sample Wilcoxon signed-rank test, we found that MutGen achieves statistically significant improvements ($\alpha$ = 0.05) in mutation scores compared with EvoSuite, EvoSuitemut, and Genvanilla. However, there is no statistically significant difference in coverage, as most subjects already achieve near 100% coverage across all approaches.
@@
In terms of overhead, MutGen consumes fewer tokens on HumanEval-Java but more tokens on Leetcode-Java. The Leetcode-Java subjects are solutions to medium- and hard-level algorithmic problems, with more complex logic than those in HumanEval-Java. As a result, PITest generates more mutants for them (19 vs. 11), and Llama-3.3 requires more test cases to kill all mutants. Compared with Genvanilla, EvoSuite, and EvoSuitemut, in Leetcode-Java, MutGen consumes more tokens while achieving higher mutation scores as well as better line and branch coverage. In contrast, on HumanEval-Java, MutGen consumes fewer tokens than EvoSuite and EvoSuitemut while still achieving superior performance in both coverage and mutation score. For the baseline approaches (i.e., EvoSuite and EvoSuitemut), a possible explanation for their higher token consumption on HumanEval-Java is that these programs contain more lines of code on average, whereas LeetCode-Java programs exhibit higher cyclomatic complexity. Longer programs tend to result in larger or more verbose test suites generated by these search-based techniques, leading to increased token usage at the test level. Interestingly, we observe an opposite trend for our LLM-based approach. On LeetCode-Java, MutGen consumes more tokens than on HumanEval-Java, despite the latter having longer programs on average. We hypothesize that this behavior is related to the use of mutation feedback in the prompt. Specifically, LeetCode-Java has a larger number of mutants per subject, leading to more uncovered or live mutants included in the mutation feedback provided to the LLM. As a consequence, the prompt becomes richer and more informative, encouraging the LLM to generate more extensive test cases to kill these mutants, which, in turn, leads to higher token consumption. However, differences of that scale in the number of tokens do not have practical implications.
@@
Overall, results suggest that MutGen significantly improves mutation scores over alternatives, with large differences, at the expense of slightly lower code coverage compared to EvoSuite for the most complex subject (Leetcode-Java). This is expected, as EvoSuite is designed to maximize coverage. For the other subject (HumanEval-Java), the reverse trend is observed. Even when EvoSuite is specifically configured as EvoSuitemut to maximize mutation scores [fraser2014large], it performs slightly worse than EvoSuite on both datasets. Therefore, to optimize mutation scores and, thus, fault detection, we recommend using MutGen, though the impact on coverage is unclear and likely negligible in practice.
Answer to RQ1 MutGen consistently achieves higher mutation scores than EvoSuite, EvoSuitemut, and Genvanilla, demonstrating its superior fault detection capability. While this may result in a slight decrease in code coverage for more complex subjects (e.g., Leetcode-Java), the difference is practically negligible.
@@
To evaluate the effectiveness and limitations of MutGen in killing mutants, we analyze the killed ratio across different mutation operators (described in [mutops]) after running MutGen for four iterations. Tables II and III present the mutant status counts for each mutation operator on HumanEval-Java and Leetcode-Java, respectively. In both tables, Column K indicates the number of killed mutants for each corresponding mutation operator. Column L shows the number of live mutants (covered by not killed), while Column NoCOV represents the number of mutants that are not covered by any test cases. Column Killed Ratio denotes the proportion of killed mutants relative to the total number of mutants.
We observe that, for frequently applied mutation operators (i.e., those appearing higher up in the tables), such as Math and NegateConditionals, MutGen is able to kill the corresponding mutants across iterations progressively. After four iterations, MutGen achieves a high killed ratio—for instance, 97.2% for Math on Leetcode-Java. The remaining 2.8% of mutants are mostly uncovered, meaning they are not exercised by any generated test cases.
In contrast, mutation operators that appear less frequently (i.e., those appearing further down in the tables) tend to exhibit lower effectiveness. For example, VoidMethodCalls shows a killed ratio of 69.2% on HumanEval-Java, and TrueReturns achieves only 57.7% on Leetcode-Java, representing the weakest effectiveness among all mutation operators in terms of killed ratio.
@@
Answer to RQ2 MutGen performs well on frequently applied mutation operators, achieving high killed ratios (e.g., 97.5% for NegateConditionals on HumanEval-Java), while operators that appear less frequently—such as VoidMethodCalls and TrueReturns—remain challenging, primarily due to dead code, uncovered statements, or subtle logic changes that the LLM fails to recognize, suggesting the need for more fine-grained semantic guidance.
@@
Based on the mutation scores obtained at the fourth iteration, code summarization appears to have the most significant impact on the effectiveness of MutGen. As discussed in Section II, code comments can sometimes mislead LLMs or omit essential details—such as input formats—that are crucial for guiding test case generation. The fixing stage also has a notable impact on effectiveness, particularly on the LeetCode-Java dataset, where many generated test cases initially fail to run. Prior work has shown that LLM-generated test cases often exhibit low execution success rates. After applying fixes, however, a substantial portion of these test cases become valid and contribute to the test objectives. Our evaluation confirms that the fixing stage significantly improves mutation scores.
Lastly, mutation feedback also enhances effectiveness. As revealed in RQ2, while many mutants generated by common mutation operators are relatively easy for LLMs to kill, mutation feedback still provides effectiveness gains by guiding the model more effectively. Future work may further investigate how mutation feedback can be tailored to help LLMs kill harder-to-detect mutants, and which mutation operators benefit the most from such guidance. We leave this exploration as a direction for future research.
Answer to RQ3 Among MUTGEN’s components, code summarization has the most significant impact on effectiveness by providing clear guidance to the LLM; the fixing stage increases the number of generated test cases that execute successfully and thereby improves the mutation score; and mutation feedback further improves performance by guiding the LLM to target harder-to-kill mutants.
@@ IV-C Threats to Validity
The threat to internal validity mainly lies in the correctness of the implementation of MutGen and the experimental scripts. To reduce this threat, we have carefully reviewed our code. Another potential threat to internal validity concerns the configuration of the employed LLM, as variations in parameters such as temperature, top-p, and maximum generation length may lead to different outcomes. In our experiments, as described in Section IV.A (Implementation), we deploy Llama-3.3 70B locally using Ollama with its default configuration and set the temperature to 0.0 to obtain more deterministic outputs, keeping all other parameters fixed across all experimental settings to mitigate this threat.
Randomness may also affect the effectiveness of MutGen. But using three repetitions in our experiments is sufficient to account for the stochastic nature of LLM outputs. Empirical observations from RQ1 support this: on HumanEval-Java, MutGen achieves average mutation scores of 89.9%, 90.2%, and 88.5% across three runs (a range of 1.7%), while on Leetcode-Java, the scores are 88.2%, 88.0%, and 91.1% (a range of 2.1%). Such differences do not affect our conclusions in any way. For the baseline generation tool, EvoSuite, we observed very similar mutation scores across three runs and all subjects in both datasets, with almost no variation. Since method-level subjects have a limited search space (around 40 lines of code per subject, as reported in Section IV.A), the evolutionary search guided by coverage converges quickly toward similar solutions across runs. Observations for EvoSuitemut, a variant of EvoSuite that maximizes mutation score, are consistent with those of EvoSuite, for the same reasons. Overall, these small differences indicate stable results across runs. Therefore, performing three repetitions is sufficient to obtain reproducible results. Reporting the average over three runs is also consistent with prior work on evaluating LLM-based approaches [yuan2024evaluating, falcao2025evaluating, gao2025trae], including test generation. This description has been clarified in Section IV.C.
@@
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [fraser2015achieving], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance relative to other LLMs is left for future work.
@@
Equivalent Mutants. Our approach uses the PITest mutation testing framework, which includes mechanisms to reduce the creation of equivalent mutants. While this mitigates the problem to some extent, the remaining equivalent mutants can reduce MutGen’s efficiency and prevent it from achieving a 100% mutation score.
Oracle Problem. We assume the original code is correct and only address assertion errors that cause Maven build failures. This simplification can lead to incorrect test oracles and false positives.
Data Leakage. While some public code in benchmark datasets may appear in the LLM’s training data (e.g., Llama-3.3), the specific mutants we generate are unlikely to be memorized by the LLM. Moreover, we compare MutGen with a vanilla prompting baseline, and the observed performance gains suggest that our improvements stem from prompt engineering and strategy design rather than memorization.
@@ V Related Work
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [chekam2017empirical, andrews2006using, inozemtseva2014coverage], and pointed in recent work from Meta [alshahwan2024automated, foster2025mutation], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [chekam2017empirical]. Although some prior work [dakhel2024effective, foster2025mutation] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work to guide LLMs in generating test cases with stronger defect-detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [fraser2011evosuite] and Randoop [pacheco2007randoop] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [fraser2015achieving] has investigated enhancing mutation scores by modifying its fitness functions, an objective closely related to ours. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [fan2023large, wang2024software, hou2024large]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [siddiq2024using, yang2024evaluation, yuan2024evaluating, zhang2024testbench, li2023nuances, pizzorno2024coverup] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [achiam2023gpt, ouyang2022training, brown2020language].
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [grattafiori2024llama, liu2024deepseek], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [ouedraogo2024large, wang2025projecttest, deljouyi2024leveraging, wang2025towards, huang2023agentcoder, kim2025llamaresttest]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [wang2025projecttest, liu2024llm, xu2025clover, tian2025fixing], Java [zhang2024testbench, wang2025projecttest, chen2024chatunitest, gu2024improving], and others [wang2025projecttest, liu2024llm, zhang2025citywalk, cheng2025rug, wang2025fine]. Other works target specific testing goals, such as detecting particular bugs [liu2024llm, foster2025mutation], API testing [cao2024automated, deng2025lrasgen, li2025llm, kim2025llamaresttest], or generating test oracles [khandaker2025augmentest, dinella2022toga, binta2024togll]. LLMs have also been integrated with traditional testing techniques, including mocking [nan2025test, zhu2025understanding], evolutionary algorithms [taherkhani2024epic], and program analysis [gu2025llm, pan2025aster, ryan2024code, pan2024multi]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [chekam2017empirical, foster2025mutation]. As discussed in Section II, high code coverage does not necessarily imply strong fault detection [chekam2017empirical]. However, as demonstrated in our experiments (Section IV-B), improving mutation score often leads to increased code coverage as a by-product.